# CayleyPy RL: Pathfinding and Reinforcement Learning on Cayley Graphs

### v1

@@ 1.3. Finite Cayley graphs: pathfinding techniques, results and open problems
Estimating diameters of Cayley graphs is a hard problem. One of its applications is latency estimation for processor interconnection problem [Akers89],[Cooperman91], [Heydemann97] - since diameter corresponds to maximal latency. It is also of interest for community of puzzles enthusiasts - like Rubik’s cube. Diameter is also called a ”God’s number” - the number of moves the worst configuration can be solved by the best (optimal) algorithm. It is widely unknown for most of the puzzles, its determination for the Rubik’s cube 3x3x3 for the standard choices of generators required significant efforts and computational resources [Rokicki14]. It is not known precisely for higher cubes, neither for 3x3x3 Rubik’s with STM, ATM metrics.A precise conjectures on 4x4x4 and 5x5x5 Rubik’s cube diameters appeared recently [Hirata24], collections of conjectures and estimates for many puzzles can be found on speedsolving website.
@@
As we will discuss below novel AI methods might provide insights via computational experiments to such kind of questions.
2. Reinforcement learning (RL) and graphs pathfiding
2.1. Reminders: RL, sparse reward problem, Bellman equation for graph distances, Q-algorithm, pathfinding
@@
RL in brief. Reinforcement learning is an important direction in AI that differs from supervised and unsupervised learning in the following ways: it focuses on optimizing cumulative rewards, and it requires the agent to, in a sense, ”create its own training set”. The agent must choose the best strategies to ”explore the space” and select training examples on its own, in contrast to supervised learning, where the training set is provided in advance.
RL ”equivalence” to graph pathfinding. The setup of reinforcement learning is very similar to the setup of graph pathfinding: states - are nodes of the graph, actions corresponds to edges/transitions to neighbor states by these edges, rewards corresponds to weights on the edges. The cumulative reward corresponds to length of paths (with respect to edges weights). Typically in graph theory shortest paths are of interest - that means ”reward” is minimized, not maximized as in RL, so it is better to use term not the ”reward”, but ”penalty”.
Unweighted graphs and sparse reward problem. The sparse reward problem is a well-known challenge in reinforcement learning (RL), and it specifically arises in pathfinding on Cayley graphs and, more generally, on any unweighted graph, which makes standard RL approaches perform poorly. The sparse reward problem occurs when rewards are given very rarely, making it difficult to generate data for training. This situation is particularly prevalent in graph pathfinding on large graphs, where the task is to find a path to a single node out of trillions of others. A reward is only given once the desired node is reached, but it is practically impossible to find it with a ”cold start” through random exploration. It’s important to note that the condition of all weights being equal to one plays a significant role. If the weights were different, one could reasonably adopt a strategy of selecting the step with the best local reward. However, when all weights are identical, there is no local preference to guide the agent in making the right move. There are several strategies to address this issue, and below, we propose a new approach.
Bellman value equation is equation for distance functions. The value function (or ”position estimation”) in RL exactly corresponds to the distance functions on graphs. To make that analogy more precise: consider an unweighted graph $G$ (all edges weights are equal to 1), choose some node $e$ (e.g. identity of a group in Cayley graph), consider a function $d(g)$ the distance of the shortest path from $g$ to $e$ (it is ”word metric” in group theory). Then there is a simple relation:
$d(g)$ $=1+\min_{n:~{}neigbours~{}of~{}g}d(n)\text{~{}~{}~{}~{}~{}~{}~{}~{}~{} Bellman equation}$ (2.1)
$d(e)$ $=0\text{~{}~{}~{}boundary condition}$ (2.2)
@@
This relation is simple, but it is fundamental - it is a particular case of the Bellman’s equation. The properties can be summarized in the following simple proposition, which is easy to see inductively considering neghbor nodes to $e$, then neighbors of neighbors and so on.
@@
For any finite connected graph $G$ the only solution to the pair of equations above are given by the distance function (length of the shortest path) to the node $e$. Moreover the only solutions to the Bellman’s equation (without boundary condition) correspond to distance to some sets of nodes in $G$ (up to adding a constant). That set can be recovered as minimums of $d(g)$.
Sense of Bellman equation - local condition for globally (cumulatively) defined object. The Bellman equation is equation for the length of the shortest path - that is globally/cumulatively defined object, but the condition itself is local - only at the neighbor of each point.
Idea of Q-learning and deep Q-learning algorithms for value/distance function. These algorithms provide a way to compute solutions of the Bellman equation, which means to find the distance functions on graphs. One starts with some initialization for $d(g)$, then compute next iteration:
$d_{k}(g)=1+\min_{n:~{}neigbours~{}of~{}g}d_{k-1}(n),d_{k}(e)=0$ by applying the right hand side of the equations 2.1. There are theoretical guarantees that such a process will converge to the solution of the Bellman equation (and boundary condition).
Idea of deep Q-learning algorithm. The deep Q-learning is used in the situations where the graph is large (can be of google size). One assumes that for each node of graph there is a feature vector attached. One considers a neural network which takes a feature vector as an input and should output the approximate solution for Bellman equations. The training process is organized as follows. Each epoch one selects some subset of nodes and from current neural network $d(g)$ computes the right hand side of the equations 2.1: $t(g)=1+\min_{n:~{}neigbours~{}of~{}g}d(n),t(e)=0$, denoting $t(g)$ as a result of that computation. Afterwards one runs gradient descent to minimize the error between $d(g)$ and $t(g)$. If the error would be zero that means that $d(g)$ solves the Bellman equations. And thus in graph theory setup it produces the distance function (”word metric” in the case of Cayley graphs). Typically the choice of the training subset of nodes is made with the help of walking over the graph combining both random steps and steps guided by the model. We will discuss that in details below.
@@
Graph path finding. To find a path from a given node $g$ to the selected node $e$, one needs to run a greedy search algorithm. Starting from node $g$, the algorithm inspects all its neighbors, evaluates the distance function for each of them, and then chooses the node with the minimum value of the distance function, moving to that node. The process is repeated until the destination node is found or the step limit is exceeded. If the distance function accurately reflects the true distance on the graph (i.e., the length of the shortest path), this algorithm will find the shortest path in an optimal way, performing the least number of operations. However, in the case of deep Q-learning, the predictions of the neural network are typically not very precise, and the greedy algorithm can get stuck in local minima. To overcome this challenge, more advanced graph search algorithms, such as beam search or A-star, are used. This will be discussed further below.
2.2. Diffusion distance pathfinding - CayleyPy-1
@@
Let us briefly remind the novel pathfinding approach proposed in our previous paper [Chervov25]. Its efficiency has been demonstrated on the Rubik’s cube where it significantly outperforms all possible competitors providing shorter solutions than any existing approach for 3x3x3,4x4x4,5x5x5 cubes. It is also more computationally efficient and in some sense it is more simple.
@@
The idea of the approach is to work not with the true distance on the graph, but with the diffusion distance which might be thought as a length of the random path. The diffusion distance is easy to estimate by the number of steps of random walks. Generation of random walks is much more fast than computations of the neural networks predictions (the only way to generate target in Q-learning), thus generation of the training data is much more efficient than in Q-learning approach, though the training data might be of weaker quality.
The assumption is that for each node of a graphs there is feature vector attached. For example for permutation groups it is a vector representing a permutation $(p_{0},p_{1},p_{2},...p_{n-1}),p:i\to p_{i}$ We also assume that there is a selected node (e.g. identity of the group) where we need to find a path from any other node.
Core of the approach consists of the three steps:
@@
    Generation of training set by random walks. Generate $N$ random walks trajectories all starting from a selected node $e$ (identity of a group). Each random walk trajectory consists of up to $K_{\text{max}}$ steps, where $N$ and $K_{\text{max}}$ are integer parameters of the method. For some nodes encountered during the random walks, we store a set of pairs $(v,k)$, where $v$ represents the vector corresponding to the node and $k$ is the number of steps required to reach it via the random walk.
@@
    Generation of training set by random walks. The generated set of pairs $(v,k)$ serves as the training set for the neural network. Specifically, $v$ serves as the ’feature vector’ (the input for the neural network), and $k$ represents the ’target’ (the output that the network needs to predict). Thus, the neural network’s predictions for a given node $v$ estimate the diffusion distance from $v$ to the selected destination node.
@@
    Graph search guided by neural network. Beam search. This step finds a path from a given node to the destination node. The neural network provides heuristics on where to make the next steps, while the graph pathfinding technique compensates for any possible incorrectness in the neural network predictions. The beam search pathfinding method is quite simple, but has proven to be the most effective for us and works as follows. Fix a positive integer $W$ a parameter known as the ”beam width” (or ”beam size”). Starting from a given node, we take all its neighboring nodes and compute the neural network predictions for all of them. We then select the $W$ nodes that are closest to the destination according to the neural network (i.e., the predictions have smaller values). For these selected $W$ nodes, we take their neighbors, drop duplicates, and again compute the neural network predictions, choosing the top $W$ nodes with the best (i.e., minimal) predictions. This process is repeated until the destination node is found (or limit of steps is exceeded).
@@
The last step is common for all methods - based on Q-learning or diffusion distance. One uses a neural network predictions as heuristic function and uses graph search algorithms to find a path. We are strongly in favor of beam search, while DeepCube team [McAleer19, Agostinelli19, Agostinelli24, Agostinelli24q] uses A-star algorithm and develops its modifications.
2.3. Beam search effectivity - toy model of research community
@@
To some extent, the success of the proposed scheme is due to beam search and our efficient implementation supporting large beams. We started with a modified greedy search, then moved to Metropolis and A-star, before finally finding relief in beam search. The larger the beam size, the more effective beam search becomes (probability to find path increases and paths become shorter). To maximize beam sizes, we developed an original and efficient PyTorch implementation capable of supporting beam sizes with millions of nodes and beyond (basic, advanced versions). On one hand, beam search is arguably one of the simplest methods. On the other hand, there are several key aspects worth highlighting.
@@
The primary reason we need beam search is to avoid the problem of local minima. Greedy search can get stuck in them, whereas beam search can bypass them entirely if the beam size is larger than the depth of the local minima. It’s similar to how an elephant wouldn’t even notice small dips in the terrain, while for an ant, they could be insurmountable obstacles.
@@
Another important aspect of beam search is that increasing the beam size not only linearly increases computation but also linearly increases memory consumption. On one hand, typically memory - not computational time - is the main bottleneck of the method. On the other hand, having the largest possible beam size is crucial for its performance. Moreover, there is no simple way to reproduce the results of beam search with a beam size of $B+1$ using any algorithm that only consumes memory equivalent to beam size $B$.
To explain this point, we use an analogy with the work of a research community.
@@
Imagine two researchers working together on a creative problem (a beam size of 2). The key factor is the exchange of ideas—if one researcher makes a breakthrough, they share it with the other, allowing both to continue their search from the breakthrough point. Without this exchange, the second researcher might get stuck in a dead end. This sharing of ideas helps avoid wasting time on unproductive directions. The mechanism described above explains why increasing the beam width is so effective—it acts as an exchange of ideas, focusing on breakthroughs rather than following unproductive paths. Essentially, this is what beam search does: it selects the top nodes across the entire beam neighborhood rather than picking the best few for each individual node. This means that if the neighbors of one node are significantly closer to the destination (analogous to a breakthrough), those nodes are retained in the beam, while the neighbors of less successful nodes are discarded. In other words, the descendants of some nodes do not survive into the next generation. A kind of natural selection process.
From the above, it is clear that the size of a research community matters. Reducing the community size by half will not simply cut the research output in half — it will decrease it exponentially. That aspect is different from the more routine work - cutting it twice output will be cut twice. The same principle applies to beam search: the beam size is crucial, and there is no simple way to achieve the performance of a beam size $B$ using a smaller beam.
@@
2.4. Novel method - combining diffusion distance with deep Q-learning
Rationale. Why we need that? The diffusion distance approach described above is quite simple and efficient however there is practical and theoretical disadvantage. The practical side consists of somewhat surprising phenomena that enlarging the training set (generating more random walks) does not always improve the quality of the model - such stagnation was observed for Rubik’s cube cases in our previous paper and it is even more prominent for LRX graphs. It is actually related to the quite clear theoretical fact: diffusion distance is not monotonically dependent on the true distance - one node can be further than another for the true distance, while oppositely ranged for the diffusion distance. (That problem is not only for random walks estimation, but for the diffusion distance itself (i.e. limit of infinite data - computed theoretically (analysis) ). So diffusion distance is easy to estimate, but it is not precise. On the other hand DQN and Bellman equation allows to approximate the true distance, but for the price of more heavy computations. The idea of the proposal below is to combine less precise but light weight diffusion distance with more heavy and more precise Bellman equation approach. And to avoid the sparse reward problem.

### v2

@@ 1.1. Broader context
Estimating diameters of Cayley graphs is a hard problem. One of its applications is latency estimation for the processor interconnection problem [Akers89],[Cooperman91], [Heydemann97] since diameter corresponds to maximal latency. It is also a question of interest for community of puzzles enthusiasts — like Rubik’s cube. In the context of puzzles the diameter is often called the ”God’s number” — the number of moves the best algorithm can solve the worst configuration. It is widely unknown for most of the puzzles. Its determination for the Rubik’s cube $3\times 3\times 3$ for the standard choices of generators required significant efforts and computational resources [Rokicki14]. It is not known precisely for higher cubes, neither for $3\times 3\times 3$ Rubik’s with STM, ATM metrics. Precise conjectures on $4\times 4\times 4$ and $5\times 5\times 5$ Rubik’s cube diameters appeared recently [Hirata24], collections of conjectures and estimates for many puzzles can be found on speedsolving website.
@@
As we will discuss below, novel AI methods might provide insights into such questions through computational experiments.
2. Reinforcement learning (RL) and graphs pathfinding
The last subsection (2.4) presents our new proposal for combining reinforcement learning (RL) training with pretraining based on the diffusion distance approach introduced in our previous work [Chervov25], which is briefly reviewed in Section 2.2. Both methods rely on beam search, recalled in Section 2.3. The first subsection (2.1) provides a brief overview of reinforcement learning and explains how it naturally fits into the graph pathfinding framework — a simple yet not thoroughly discussed perspective in the existing literature, to the best of our knowledge.
2.1. RL setup for pathfinding and Bellman equation
@@
Reinforcement learning is an important subfield of AI, which is different from supervised and unsupervised learning in the following respects: it aims to optimize cumulative rewards, and it requires the agent to, in a sense, ”create its own training set”. The agent should choose the best strategies to ”explore the space” and select training examples by himself, in contrast to supervised learning, where the training set is given a priori. Here we will briefly discuss it from the graph theory perspective.
Bellman equation and shortest paths. Let $d(s)$ be the distance function — the length of the shortest path from the state $s$ to the target node $e$. Denote the set of all neighbors of $s$ as $N(s)$. It is obvious that, the distance function $d(s)$ satisfies the Bellman equation:
$d(s)=1+\min\limits_{t\in N(s)}d(t)$ (2)
and the boundary condition
$d(e)=0.$ (3)
The equation is elementary but plays an important role in reinforcement learning and its deep learning generalizations discussed below.
@@
The Bellman equation implies a simple proposition, which is easy to see inductively considering neighbor nodes to $e$, then neighbors of neighbors, etc. (actually, this is how the breadth-first search algorithm finds the shortest path).
@@
For any finite connected graph $G$, the only solution to the equations (2) and (3) is given by the distance function $d(s)$. Moreover, the only solutions to Bellman’s equation (2) (without boundary condition) correspond to the distance to some sets of nodes in $G$ (up to adding a constant). That set can be recovered as minimums of $d(s)$.
RL ”equivalence” to graph pathfinding. Here we remind the standard form of Bellman equation for the value function $v_{*}(s)$ ([Sutton18, p.62]) and relate it to the form given above. Taking any action $a$ in state $s$ the agent receives a reward (penalty) $r(s,a)$, and denote discount factor to be $\gamma\in(0,1]$, then Bellman optimality equation:
$v_{*}(s)=\max\limits_{a}\big\{r(s,a)+\gamma\mathbb{E}_{s^{\prime}}v_{*}(s^{\prime})\big\}$
Considering environment to be deterministic, $\mathbb{E}_{s^{\prime}}v_{*}(s^{\prime})=v_{*}(a\circ s)$ and taking r(s,a)=1r(s,a)=-1, one gets:
$v_{*}(s)=-1+\gamma\max\limits_{a}\big\{v_{*}(a\circ s)\big\}.$ (4)
Now observe that (2) turns into (4) if $\gamma=1$ and $d(s)=-v_{*}(s)$. By definition of the value function $v_{*}(e)=0$ which is the same as (3).
Solving Bellman - simple iteration method (dynamic programming). Bellman equation on unknown function $d$ has the form $d=F(d)$, thus to solve it one can use method of simple iteration: $d_{k+1}=F(d_{k})$. Thus, solving Bellman equation is conceptually rather simple - starting with some initialization $d_{0}(s)$, the new approximation for the distance function is iteratively updated by by the right hand side of the Bellman equation (method is also called dynamic programming in RL context), i.e.:
$d_{k}(s)=1+\min_{t\in N(s)}d_{k-1}(t),\quad d_{k}(e)=0.$
Algorithm 1 Simple iteration (Dynamic programming)
1:Hyperparameters: $\epsilon>0$ (tolerance), $\alpha\in(0,1]$ (learning rate), NN (maximum iterations)
2: Initialize $d_{0}(s)$ for all states $s\in\mathcal{S}$ (e.g., by zeros)
3:for $i=0$ to NN do
4:  $d_{i+1}(e)\leftarrow 0$
5:  for each state $s\neq e$ do
6:   $d_{i+1}(s)\leftarrow\alpha(1+\min\limits_{a}d_{i}(a\circ s))+(1-\alpha)d_{i}(s)$
7:  end for
8:  $c_{i}\leftarrow\mathrm{corr}(\boldsymbol{d}_{i},\boldsymbol{d})$
9:  if $\|\boldsymbol{d}_{i}-\boldsymbol{d}\|<\epsilon$ then
10:   break
11:  end if
12:end for
13:Output: Number of iterations $i$, vector of correlations $\boldsymbol{c}$
Examples of solving Bellman equation. Initialization dependence. Let us illustrate solving Bellman equation on Cayley graphs and demonstrate some unexpected dependence on initializations. We observe that if initialization is positive convergence takes not more than diameter steps (it is no longer true for initialization with negative values). However convergence might not be monotonic - depending on initialization, it might happen that first solution deviates from the final result, and only after several steps, begins to converge again.
For small graphs such as Pocket Cube or LRX with $n\leqslant 10$ the true distances $d(s)$ can be easily calculated via BFS for each state. Therefore, we can use these true distances as stopping criteria: when $\|d_{n}(s)-d(s)\|<\varepsilon$, we are finished. In addition, we measure performance on each iteration by calculating the correlations between $d(s)$ and $d_{k}(s)$.
The main question of interest here was how the initialization influences the performance. That is important to understand in view of our proposal to use initialization coming from the diffusion distance. We tried different initializations $d_{0}(s)$:
  • zero initialization $d_{0}(s)=0$ for all $s$
  • initialize each of $m$ layers by a specific value (e.g., first $k$ layers by true distances, last mkm-k layers by zeros, $1\leqslant k<m$)
  • Manhattan distance to the target: $d_{0}(s)=\sum|s_{k}-k|$
  • Hamming distance to the target: $d_{0}(s)=|\{k\colon s_{k}\neq k\}|$
  • random integers form 0 to $m$ or random values from $\mathcal{N}(0,\sigma^{2})$
  • outputs of a model predicting diffusion distances (see next subsection or formula (10))
Figure 3. Pearson correlation between $d_{i}(s)$ and $d(s)$ on each iteration of the DP algorithm, $\epsilon=10^{-3}$ (Pocket Cube)
These experiments revealed interesting phenomena presented in figure 3. For some specific initialization the performance first degrades, and only after that achieves the convergence to the true distance which is guaranteed by the general theory. The number of steps for convergence does not exceed the diameter of the graph if initialization is positive, however, it can be larger in presence of negative values. Initialization coming from the machine learning model (”catboost”) trained on diffusion distance (same as ”warm-up step” in our proposal) improves the convergence, though not in a radical manner.
Deep Q-learning algorithm. Tabular methods like DP or Q-learning are not applicable in situations where the graph is too large. A family of deep Q-learning methods like [Mnih13] emerged to resolve this issue. Typically, these algorithms leverage a neural network to predict targets like right-hand sides of 4 or 2:
$f_{\boldsymbol{\theta}}(s)=r+\max_{a}Q_{\boldsymbol{\theta}}(s,a).$
The algorithm 2 is a deep Q-learning method designed to solve the Bellman equation 2 (here we assume that $s$ is an embedding of a graph node suitable for input of the neural network $f_{\boldsymbol{\theta}}(s)$). Typically, the choice of the training subset of nodes is made by walking over the graph, combining both random steps and steps guided by the model. After the training is completed, the output of the model $f_{\boldsymbol{\theta}}(s)$ should hopefully serve as a good approximation to the true distance from $s$ to $e$.
Algorithm 2 Deep Q-learning
1:Hyperparameters: NN (number of epochs), $M$ (subset size)
2: Initialize weights $\boldsymbol{\theta}$ of the neural network $f_{\boldsymbol{\theta}}(s)$
3:for $i=0$ to NN do
4:  select a random subset of nodes $S_{i}$, $|S_{i}|=M$, $e\notin S_{i}$
5:  for each state $s\in S_{i}$ do
6:   $d_{i}(s)\leftarrow 1+\min\limits_{t\in N(s)}f_{\boldsymbol{\theta}}(t)$
7:  end for
8:  calculate MSE loss $\mathcal{L}(\boldsymbol{\theta})=\frac{1}{M}\sum\limits_{s\in S_{i}}\big(d_{i}(s)-f_{\boldsymbol{\theta}}(s)\big)^{2}$
9:  calculate $\nabla_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta})$ and do the gradient step
10:end for
@@
Graph path finding. To find a path from a given node $s$ to the target $e$, one needs to run a greedy search algorithm, which, starting from node $s$, inspects all its neighbors and evaluates distance function for them, then chooses the node with the minimum value of the distance function and makes a move to that node. The process is repeated until the number of the destination node found at the limit of the step is exceeded. If the distance function is the true distance on the graph (the shortest path length), such an algorithm finds the shortest path optimally (least possible operations). However, the predictions of the neural network $f_{\boldsymbol{\theta}}(s)$ from the deep Q-learning algorithm 2, are typically not very precise, and the greedy algorithm will be stuck in local minima. To overcome that difficulty, one used more advanced graph search algorithms like beam search, A-star, etc.
A guided beam search approach is described in algorithm 3; on each iteration it maintains $W$ states with lowest estimations to the target state. If algorithm 3 can find the path to the target in NN iterations, it returns the length of this path.
2.2. Diffusion distance pathfinding (CayleyPy-1)
@@
Let us briefly remind the alternative pathfinding approach proposed in our previous paper [Chervov25]. Its efficiency has been demonstrated on the Rubik’s cube, where significantly outperforms all possible competitors, providing shorter solutions than any existing approach for $3\times 3\times 3$, $4\times 4\times 4$, $5\times 5\times 5$ cubes. In some sense, it is simpler and more computationally efficient.
Algorithm 3 Guided beam search pathfinding
1:Hyperparameters: $s_{0}$ (starting state), $M$ (step limit), $f(s)$ (distance estimator), $N(s)$ (neighbors of $s$), $W$ (beam width)
[…]
3:for $i=0$ to $M$ do
4:  $B^{\prime}\leftarrow\varnothing$
5:  for each state $s\in B$ do
6:   if $s=e$ then
7:    return $i$
8:   end if
9:   $B^{\prime}\leftarrow B^{\prime}\cup N(s)$
10:  end for
11:  $B^{\prime}\leftarrow B^{\prime}\setminus B$
12:  sort states $s\in B^{\prime}$ by the distance estimations:
$f(s_{(1)})\leq f(s_{(2)})\leq\ldots\leq f(s_{(|B^{\prime}|)})$
13:  $B\leftarrow\{s_{(1)},\ldots,s_{(\min\{W,|B^{\prime}|\})}\}$
14:end for
The idea is to work not with the true distance on the graph but with the diffusion distance, which is the length of the random path. The rigorous definition of the diffusion distance is given by the formula (10) in the supplementary section.
@@
The diffusion distance is easy to estimate by the number of steps of random walks. Generating random walks is much faster than computing neural network predictions. Thus, the training data generation is much more efficient than in Q-learning approaches, though the training data might be of poorer quality.
The core of our approach consists of three steps:
@@
    Generation of the training set via random walks. Generate $N$ random walk trajectories starting from the selected node $e$. Each random walk trajectory consists of up to $K_{\text{max}}$ steps, where $N$ and $K_{\text{max}}$ are integer parameters of the method. If a random walk visits node $s$ for the first time at step $k$, we store the pair $(s,k)$ in our training set.
@@
    Training a model. The generated set of pairs $(s,k)$ serves as the training data for a supervised machine learning model $f(s)$, typically a neural network. For a given node $s$ this model predicts the estimated diffusion distance from $s$ to the target $e$.
@@
    Guided beam search. Here we apply the algorithm 3 to find a path from a given state $s$ to the target node $e$. The model $f(s)$ trained on the previous step provides heuristics on where to make the next steps, while the beam search technique compensates for any possible incorrectness in the predictions of the neural network. Having fixed a positive integer $W$ (”beam width” or ”beam size”), on each step we select $W$ nodes that are closest to the destination according to the neural network. Then we check $f(s)$ for all neighbors of these best $W$ states, drop duplicates, and repeat the process until the destination node is found or the step limit is exceeded.
@@
The last step is common for any method, whether it is based on Q-learning or diffusion distance. One uses a neural network predictions as a heuristic function and graph search algorithms to find a path. We are strongly in favor of beam search since it is quite simple, but has proven to be the most effective for us, while the DeepCube team [McAleer19, Agostinelli19, Agostinelli24, Agostinelli24q] uses the A-star algorithm and develops its modifications.
A simple but useful modification of the diffusion distance method leverages non-backtracking which is applied during steps (1) and (3). On both occasions, we store the history states visited in the $m$ previous steps to avoid returning to them. The history of non-backtracking random walks contains $O(mN)$ states, and all trajectories are not allowed to visit any of these states and their neighbors in the next step. For non-backtracking beam search we do not include the states from the history on line 9 of the Algorithm 3.
2.3. Beam search effectivity. Toy model of research community
@@
To some extent, the success of the proposed above is due to beam search and our efficient implementation supporting large beams. We started with a modified greedy search, then moved to Metropolis and A-star, before finally finding relief in beam search. The larger the beam size, the more effective beam search becomes (probability to find path increases and paths become shorter). To maximize beam sizes, we developed an original and efficient PyTorch implementation capable of supporting beam sizes with millions of nodes and beyond (basic, advanced versions). On one hand, beam search is arguably one of the simplest methods. On the other hand, there are several key aspects worth highlighting.
@@
The primary reason we need beam search is to avoid the problem of local minima. Greedy search can get stuck in them, whereas beam search can bypass them entirely if the beam size is larger than the depth of the local minima. It is similar to how an elephant wouldn’t even notice small dips in the terrain, while for an ants, they could be an insurmountable obstacles.
@@
Another important aspect of beam search is that increasing the beam size increases both computation and memory consumption linearly. Typically, memory – not computation time – is the main bottleneck. However, maximizing the beam size is crucial for performance. Moreover, there is no simple way to reproduce the results of beam search with a beam size of $B+1$ using any algorithm that only consumes memory equivalent to beam size $B$.
@@
To explain this point, we will use the analogy with the work of the research community. Imagine two researchers working together on a creative problem (a beam size of 2). The key point is the exchange of ideas: if one researcher makes a breakthrough, they share it with the other, allowing both to continue their search from the breakthrough point. Without this exchange, the second researcher might get stuck in a dead end. This sharing of ideas helps avoid wasting time in unproductive directions. The mechanism described above explains why increasing the beam width is so effective: it acts as an exchange of ideas, focusing on breakthroughs rather than following unproductive paths. Essentially, this is what beam search does: it selects the top nodes across the entire beam neighborhood rather than picking the best few for each individual node. This means that if the neighbors of one node are significantly closer to the destination (analogous to a breakthrough), those nodes are retained in the beam, while the neighbors of less successful nodes are discarded. In other words, the descendants of some nodes do not survive into the next generation. A kind of natural selection process.
These observations show that community size has a disproportionately large impact on research productivity. Halving the number of researchers does not simply halve the output — it causes a super-linear (often exponential) decline. In contrast, for routine tasks, a 50% reduction in resources usually yields about a 50% reduction in output. The same principle applies to beam search: the beam size is crucial, and there is no simple way to achieve the performance of a beam size $B$ using a smaller beam.
@@
2.4. Combining diffusion distance with deep Q-learning
The diffusion distance approach described above is quite simple and efficient; however, there are practical and theoretical disadvantages. The practical side consists of somewhat surprising phenomena that enlarging the training set (generating more random walks) does not always improve the quality of the model; such stagnation was observed for Rubik’s cube cases in our previous paper, and it is even more prominent for LRX graphs. It is actually related to the quite clear theoretical fact: diffusion distance is not monotonic function of the true distance. This problem is not only for random walks estimation, but also for the diffusion distance itself (e.g., see computed theoretically analysis). So, diffusion distance is easy to estimate, but it is not precise. On the other hand, DQN and Bellman’s equation allow us to approximate the true distance, but for the price of more heavy computations. The idea of the proposal below merges lightweight diffusion distance with a heavier and more precise Bellman equation approach, and at avoiding the sparse reward problem.